# Bifurcations in the Kuramoto model with external forcing and higher-order interactions

### v1

@@ III.4 Evolution of Bifurcation curves
As $K_{1}$ and $K_{23}$ varies, the bifurcation manifolds change. For the sake of simplicity we shall refer to the two saddle-node branches as the green (thick green line) and black (thin black line) branches, referring to the colors displayed in Fig. 1. Figure 3 shows saddle-node and Hopf bifurcation curves in the $F\times\Omega$ plane for a few selected values of $K_{1}$ and $K_{23}$. For $K=1$, panels (a)-(c), the green branch of the saddle-node manifold is almost independent of $K_{23}$, whereas the black branch, created for $K_{23}\approx 5.83$ (see below), grows as $K_{23}$ increases. For $K_{23}=5.8$, panel (a), the black branch has collapsed into the green branch, but its Takens-Bogdanov point remains with its associated Hopf curve. Panels (d)-(f) have $K_{23}=7$ and increasing values of $K_{1}$. The green branch now shrinks and disappears at $K_{1}=2$. For $K_{1}>2$ only one branch and its attached Hopf curve survive, similar to the case $K_{23}=0$.
@@
Figure 3: Saddle-node (solid) and Hopf (dashed) curves for different values of $K_{1}$ and $K_{23}$. In panels (a)-(c) $K_{1}=1$ and $K_{23}$ is 5.80, 5.93 and 6.5 respectively. Notice that the thick green saddle-node branch and its correspondent Hopf curve are almost independent of $K_{23}$, whereas the other ones grow as $K_{23}$ increases. In (d)-(f) $K_{23}=7$ and $K_{1}$ is 1.5, 1.95 and 2.1 respectively. The thick branch disappears at $K_{1}=2$. Black squares show Takens-Bogdanov points.
@@
More information about the critical values described above can be obtained by setting $F=\Omega=0$ in Eqs. (15) and (16). This amounts to set $A=0$, Eq.(18). Solving for $K_{23}$ we find $K_{23}=(a-K_{1})/r^{2}$. Imposing $\partial K_{23}/\partial r=0$ at $r=r_{c}$ we obtain
@@
For $K_{1}=1$ we find $r_{c}=\sqrt{2}-1$ and $K_{23}(r_{c})\approx 5.83$. Below this point the two legs of the black branch collapse into the lower end of the green branch (see Fig. 3(a)). Notice that $K_{23}(r_{c})$ is larger than $K_{23}^{*}\approx 5.306$, which means that two Hopf bifurcation curves exists attached to a single SN tongue. This phenomena will occur for $K_{1}=1$ as long as $K_{23}$ is lying in [5.306, 5.83] interval. Decreasing $K_{23}$ even more would show the two Takens-Bogdanov points coalescing at $K_{23}^{*}$ and the Hopf curves disappearing.
@@ IV Forcing parameters as fixed
The authors of [40] found a pitchfork bifurcation for $K_{1}=2$, independent of $K_{23}$, and a saddle-node bifurcation along a certain curve in the $K_{1}\times K_{23}$ plane. It turns out that the situation in the presence of forcing is quite different: for any finite value of the forcing parameters, the pitchfork bifurcation is replaced by another saddle-node curve; moreover, a manifold of Hopf bifurcations also appears.
@@
and its coordinates in the $K_{1}\times K_{23}$ plane are given by (32) and (33) computed at $r_{\rm TB}$.
@@
In Figure 4 we show the bifurcation curves in the $K_{1}\times K_{23}$ plane, for a few values of $F$ and $\Omega$. In panels (a)-(e), $\Omega=0.5$ and $F$ takes the values $0.45$, $0.47$, $0.49$, $0.499$ and $0.59$, respectively. Initially, we have the saddle-node branches (thin black and thick green solid lines) in some sense connected by the Hopf branches (red and blue dashed lines). The point where the Hopf curves join is $K_{23}^{*}$, as discussed in section III.2. As $F$ grows, the lower SN branch grows and crosses the upper one, which slowly receed to negative values of $K_{1}$. When $F\to\Omega$, the TB points slide off to infinity. In panel (f) we show that for small values of $F$ and $\Omega$, the bifurcation diagram from [40] is recovered.
Figure 4: Saddle-node and Hopf curves for different values of $F$ and $\Omega$. In panels (a)-(e) $\Omega=0.5$ and $F$ is 0.45, 0.47, 0.49, 0.499 and 0.59, respectively. In panel (f), $\Omega=0.01$ and $F=0.02$. Black squares show Takens-Bogdanov points.
Figure 5: Heat map of the time averaged order parameter $\langle r\rangle$ showing the phase transitions promoted by the bifurcation lines, for $\Omega=0.5$ and $F=0.45$, as in fig. 4a.
Figure 5 shows a heat map of the time averaged order parameter $\langle r\rangle$, along with the bifurcation lines. The transitions in this case are not as abrupt as they appear in Figure 1-(b), with the exception of one of them, in the upper branch of the saddle-node manifold. The Hopf bifurcations, on the other hand, are not very noticeable, because in this case what happens is that a stable fixed point with $\langle r\rangle>0$ becomes unstable and a periodic orbit is created which has approximately the same $\langle r\rangle$.
@@ V Conclusions
In this work we extended the bifurcation analysis of the forced Kuramoto model [14] to the case of asymmetric higher order interactions, as discussed in [40]. We have described in detail how these type of higher order interactions impact the bifurcation diagram obtained originally in [14] in the $F\times\Omega$ plane and also how the intensity and frequency of the external force change the diagram described in [40] in the $K_{1}\times K_{23}$ plane.
We have shown that two branches of the saddle-node (SN) bifurcation manifold can co-exist for some values of the parameters, reflecting the bi-stability of the unforced case. From each branch a line of Hopf bifurcation issues from the Takens-Bogdanov (TB) point, dividing the space of parameters in many regions of quite distinct dynamical character, as illustrated in Figs. 1-4.
We did not compute the manifold of homoclinic bifurcations that starts at the TB point and ends at the lower end of the SN tongues [14]. There is no known algebraic procedure to compute these manifolds, i.e. they must be found by numerically scanning the periodic orbits in the vicinity of the cusp of the SN curves, requiring intense numerical calculations.
We note that higher order interactions can be defined in different ways [39, 41, 40] and most of them are not amenable to exact treatment under the Ott-Antonsen ansatz. Symmetric interactions, for instance, can lead to multi-stability [39, 41] which could lead to a proliferation of branches in the bifurcation manifolds. We leave the exploration of these cases for future work. Finally, we emphasize that the importance of understanding forced systems of oscillators goes much beyond pure mathematical interest, and has applications in biological systems such as neurons subjected to visual stimuli, artificial pacemakers for the heart and the circadian rhythms promoted by the day-night cycle.
@@ Acknowledgements.
This work was partly supported by FAPESP, grants 2021/14335-0 (MAMA), 2023/03917-4 (GSC) and 2023/15644-2 (MN), and also by CNPq, grants 301082/2019-7 (MAMA) and 304986/2022-4 (MN).

### v2

@@ III.5 Evolution of Bifurcation curves
As $K_{1}$ and $K_{23}$ varies, the projection of the bifurcation manifolds on the $(F,\Omega)$ plane changes. Figure 3 illustrates how SN and Hopf bifurcation curves change for a few selected values of $K_{1}$ and $K_{23}$. For the sake of simplicity we shall refer to the two SN branches as the green (thick green line) and black (thin black line) branches, referring to the colors displayed in Fig. 1. Also, for the sake of clarity and lack of resolution in the figures, we do not display the homoclinic bifurcation curves. For $K_{1}=1$, panels (a)-(c), the green branch of the SN manifold is almost independent of $K_{23}$, whereas the black branch, created at $K_{23}\approx 5.83$ (see below), grows as $K_{23}$ increases. For $K_{23}=5.8$, panel (a), the black branch has collapsed into the green branch, but its Takens-Bogdanov point remains with its associated Hopf curve. Panels (d)-(e) have $K_{23}=7$ and increasing values of $K_{1}$. The green branch now shrinks and disappears at $K_{1}=2$. For $K_{1}>2$ only one branch and its attached Hopf curve survive, similar to the case $K_{23}=0$. Two new regions were identified with respect to the classification in Figs. 1 and 2. Region $D^{*}$ differs from $D$ by having the only stable fixed point close to the origin, indicating disordered motion. Figure 3(f) shows the vector field in the novel region K that appears in panels 3(a) and (b):
Region K - Bi-stability between disordered motion and forced entrainment, similar to region $E$, but without the unstable and saddle fixed points observed at the left of Fig. 2(e), that collide on the SN bifurcation.
@@
Figure 3: Saddle-node (solid) and Hopf (dashed) curves for different values of $K_{1}$ and $K_{23}$. In panels (a)-(c) $K_{1}=1$ and $K_{23}$ is 5.80, 5.93 and 6.5 respectively. Notice that the thick green saddle-node branch and its correspondent Hopf curve are almost independent of $K_{23}$, whereas the other ones grow as $K_{23}$ increases. In (d)-(e) $K_{23}=7$ and $K_{1}$ is 1.5 and 2.1 respectively. The thick green branch disappears at $K_{1}=2$. Black squares show Takens-Bogdanov points. Panel (f) shows the vector field and fixed points found in Region $K$ ($F=0.15$, $\Omega=0.1$, $K_{1}=1$, $K_{23}=5.93$).
@@
More information about the critical values described above can be obtained by setting $F=\Omega=0$ in Eqs. (15) and (16). This amounts to set $A=0$, Eq.(18). Solving for $K_{23}$ we find $K_{23}=(m-K_{1})/r^{2}$, where $m=2/(1-r^{2})$. As a function of $r$, $K_{23}$ is U-shaped in the interval $(0,1)$. Imposing $\partial K_{23}/\partial r=0$ at $r=r_{c}$ we obtain
@@
For $K_{1}=1$ we find $r_{c}=\sqrt{2}-1$ and $K_{23}(r_{c})\approx 5.83$. Below this point the two legs of the black branch collapse into the lower end of the green branch (see Fig. 3(a)). Notice that $K_{23}(r_{c})$ is larger than $K_{23}^{*}\approx 5.306$, which means that two Hopf bifurcation curves exist attached to a single SN tongue. This phenomena will occur for $K_{1}=1$ as long as $K_{23}$ is in the [5.306, 5.83] interval. Decreasing $K_{23}$ even more would show the two Takens-Bogdanov points coalescing at $K_{23}^{*}$ and the Hopf curves disappearing.
@@ IV Forcing parameters as fixed
The authors of [46] found a pitchfork bifurcation for $K_{1}=2$, independent of $K_{23}$, and a saddle-node bifurcation along a certain curve in the $(K_{1},K_{23})$ plane. It turns out that the situation in the presence of forcing is quite different: for any finite value of the forcing parameters, the pitchfork bifurcation is replaced by a saddle-node curve, together with Hopf and homoclinic bifurcation lines; moreover, this structure may appear duplicated.
@@
and its coordinates in the $(K_{1},K_{23})$ plane are given by (32) and (33) computed at $r_{\rm TB}$.
Figure 4: Saddle-node and Hopf curves for different values of $F$ and $\Omega$. In panels (a)-(e) $\Omega=0.5$ and $F$ is 0.45, 0.47, 0.49, 0.499 and 0.59, respectively. In panel (f), $\Omega=0.01$ and $F=0.02$. Black squares show Takens-Bogdanov points and triangles are Bautin points, where sub and super-critical Hopf bifurcation curves meet (see Fig. 5(a)). Regions are labeled according to Figs. 1, 2 and 3.
Figure 5: (a) Zoom of Fig. 4(a) showing the joining of sub-critical (blue) and super-critical (red) Hopf bifurcation manifolds at the Bautin point (triangle). A line of saddle-node bifurcation of cycles (purple) starts from the same point. (b) Equilibrium values of the order parameter $r$ as function of $K_{1}$ for $K_{23}=4$. Solid lines for stable equilibria, dashed lines for unstable ones. Black lines display a pitchfork bifurcation for $F=0$. For $F=0.02$ and $\Omega=0.01$, red lines, the pitchfork unfolds into saddle-node bifurcations. Values of $r<0$ are shown for the sake of visualization.
@@
Figure 4 shows the bifurcation curves in the $(K_{1},K_{23})$ plane, for a few values of $F$ and $\Omega$. In panels (a)-(e), $\Omega=0.5$ and $F$ takes the values $0.45$, $0.47$, $0.49$, $0.499$ and $0.59$, respectively. Panel (a) shows the two saddle-node branches (thin black and thick green solid lines) connected by the red (super-critical) and blue (critical) Hopf branches. The point where the Hopf curves join is $K_{23}^{*}$, as discussed in section III.2, and characterize a Generalized Hopf, or Bautin, bifurcation, where a line of SN of limit cycles also issues [57], as shown in Fig. 5(a).
As $F$ grows, the lower SN branch grows and crosses the upper one, which slowly recedes to negative values of $K_{1}$. When $F\to\Omega$, the TB points slide off to infinity. In panel 4(f) we show that for small values of $F$ and $\Omega$, the bifurcation diagram from [46] is recovered. However, for any $F>0$ the pitchfork bifurcation unfolds into SN curves, as shown if Fig. 5(b), where we plot all equilibrium values of $r$ that exist for $K_{23}=4$ as functions of $K_{1}$. As a consequence, $r=0$ is no longer an equilibrium.
@@ V Conclusions
Coupled oscillators are often subjected to external forcing, introducing a competition between spontaneous and forced synchronization. Examples include artificial pacemaker devices implanted on the heart [17], responses of the visual cortex to stimuli [19] and epileptic seizures induced by flashing lights [16]. The importance of understanding forced systems of oscillators, therefore, goes much beyond pure mathematical interest. Child and Strogatz [13] used the Kuramoto model to study the effects of a periodic force. They showed that the asymptotic states depended on the force intensity $F$ and on the mismatch between the external frequency and average natural frequency of the oscillators, $\Omega$. For fixed $\Omega$ and small $F$, spontaneous synchronization is observed, whereas for sufficiently large $F$ the oscillators synchronize with the external force. More importantly, the transition between the two regimes displayed a rich set of bifurcations with periodic cycles and equilibrium points that depended very sensitively on the values of $F$, $\Omega$ and the coupling parameter $K$.
In this work we extended the bifurcation analysis of the forced Kuramoto model to include higher order interactions between the oscillators. The effects of higher order terms in states of the Kuramoto model was discussed by Skardal and Arenas, who showed that the contribution of third and forth order interactions could be put together in terms of a single parameter, $K_{23}$, but did not considered external forces [46]. Here we described in detail how these interactions impact the bifurcation diagram obtained originally in [13] in the $(F,\Omega)$ plane and also how the intensity and frequency of the external force change the diagram described in [46] in the $(K_{1},K_{23})$ plane.
We have shown that two branches of the saddle-node (SN) bifurcation manifold can co-exist for some values of the parameters, reflecting the bi-stability found in [45] for the unforced case. Lines of Hopf and homoclinic bifurcations issue from Takens-Bogdanov points, dividing the space of parameters in many regions of quite distinct dynamical character, as illustrated in Figs. 1-4. The two branches of Hopf bifurcations meet at the so called generalized Hopf or Bautin bifurcation, point where a line of saddle-node of cycles also starts. The combined effects of higher order interactions and external periodic force increases dramatically the complexity of the bifurcation diagrams, that is divided into 11 different topological regions, making the determination of the asymptotic state of the system very sensitive to parameter values.
Finally, we note that higher order interactions, as derived from phase reduction methods [52, 53], can have different functional forms, and a study of their combined effects has not yet been investigated in detail. Instead, the effect of specific terms have been studied separately, revealing distinct patterns and states. Asymmetric interaction terms are amenable to exact treatment under the Ott-Antonsen ansatz, and have been analysed in refs. [45, 29] and here. The Kuramoto model with symmetric higher order terms, on the other hand, exhibits multi-stability [26, 45, 47, 46] and anomalous transitions to synchrony [30]. In this case, the effects of external forcing might lead to a proliferation of branches in the bifurcation manifolds. We leave the exploration of these cases for future work.
@@ Acknowledgements.
We would like to thank the anonymous referees for their valuable remarks. This work was partly supported by FAPESP, grants 2021/14335-0 (MAMA), 2023/03917-4 (GSC) and 2023/15644-2 (MN), and also by CNPq, grants 301082/2019-7 (MAMA) and 304986/2022-4 (MN).